s l 0 if you drew one card from a standard deck, would it be signiicant to draw a 57 Answer the question, considering an event to beignificant. İf itprobability ibu han-equal noos No
Answer #1

No ; as all events are equally likely and probability of drawing a 5 =4/52 =0.077 (as there are 4 cards of 5)